# No-go theorem for probabilistic one-way secret-key distillation

### v1

@@ Appendix D Convexity and openness of the set of super two-extendible states
$\operatorname{supp}\!\left(\operatorname{Tr}_{B}\!\left[\overline{\rho}_{ABE}\right]\right)\subseteq\operatorname{supp}\!\left(\rho_{AB}\right),$ (67)
where $E$ is isomorphic to $B$. Note that the existence of such an extension is guaranteed by the fact that $\rho_{AB}$ is in the set of super two-extendible states. Similarly, let $\overline{\tau}_{ABE}$ be an extension of $\tau_{AB}$ such that the following containment holds:
$\operatorname{supp}\!\left(\operatorname{Tr}_{B}\!\left[\overline{\tau}_{ABE}\right]\right)\subseteq\operatorname{supp}\!\left(\tau_{AB}\right).$ (68)
@@
$\omega_{ABE}\coloneqq p~{}\overline{\rho}_{ABE}+(1-p)\overline{\tau}_{ABE},$ (70)
@@
$\operatorname{Tr}_{E}\!\left[\omega_{ABE}\right]=p~{}\rho_{AB}+(1-p)\tau_{AB}=\sigma_{AB}.$ (71)
@@
$\operatorname{supp}\!\left(\operatorname{Tr}_{B}\!\left[\omega_{ABE}\right]\right)$
$=\operatorname{supp}\!\left(\operatorname{Tr}_{B}\!\left[\overline{\rho}_{ABE}\right]\right)\cup\operatorname{supp}\!\left(\operatorname{Tr}_{B}\!\left[\overline{\tau}_{ABE}\right]\right)$ (73)
@@
Since the support of $\operatorname{Tr}_{B}\!\left[\omega_{ABE}\right]$ is in the support of $\operatorname{Tr}_{E}\!\left[\omega_{ABE}\right]$, we conclude that $\operatorname{Tr}_{E}\!\left[\omega_{ABE}\right]=\sigma_{AB}$ is a super two-extendible state for all $p\in(0,1)$ as well. Therefore, the state $\sigma_{AB}$, which is a convex combination of two arbitrary super two-extendible states, is super two-extendible for all $p\in[0,1]$, justifying that the set of super two-extendible states is convex.
To show that the set of super two-extendible states is open, let us consider the example of erased states defined in (14). The erased state $\widetilde{\Phi}^{p,k}_{AB}$ is a special case of an erased private state $\widetilde{\eta}^{p,k}_{AB}$. Appendix G establishes that an erased private state is super two-extendible for every $p\in[0,1)$. This implies that an erased state $\widetilde{\Phi}^{p,k}_{AB}$ is also super two-extendible for all $p\in[0,1)$. However, for $p=1$, $\widetilde{\Phi}^{p,k}_{AB}$ is a maximally entangled state of Schmidt rank $k$, which is not super two-extendible. In fact, this state has a min-unextendible entanglement of $\log_{2}k$. Hence, the set of super two-extendible states is open.
@@ Appendix E Proof of Proposition 1
We now establish the opposite implication. Consider an arbitrary quantum state $\rho_{AB}$ such that $E^{u}_{\min}\!\left(\rho_{AB}\right)=0$. This is true only if there exists a quantum state $\sigma_{AB}\in\mathcal{F}\!\left(\rho_{AB}\right)$ such that $\operatorname{Tr}\!\left[\Pi^{\rho}_{AB}\sigma_{AB}\right]=1$, which follows from the definition of the min-unextendible entanglement of a bipartite state. From the gentle measurement lemma [37, Lemma 5], it is known for a state $\tau$ and a projector $P$ that
@@ Appendix F Proof of Theorem 1
00 0 $\leq E^{u}_{\min}\!\left(q[1]_{X_{A}}\otimes\gamma^{k^{\prime}}_{A^{\prime}B^{\prime}}+(1-q)[0]_{X_{A}}\otimes\frac{I_{A^{\prime}B^{\prime}}}{d_{A^{\prime}}d_{B^{\prime}}}\right)$ (85)
@@
Inspecting (85) and (90), it follows that $E^{u}_{\min}\!\left(\eta^{q,k^{\prime}}_{A^{\prime}B^{\prime}}\right)=0$. By applying Lemma 1, using the fact that $-\frac{1}{2}\log_{2}\!\left(\frac{q}{k^{\prime 2}}+1-q\right)\geq 0$ for all $q\in[0,1]$ and integer $k^{\prime}\geq 2$, and solving the equation $0=-\frac{1}{2}\log_{2}\!\left(\frac{q}{k^{\prime 2}}+1-q\right)$, along with the assumption that $k^{\prime}\geq 2$, we conclude that $q=0$. As such, the expected number of secret key bits distilled by the channel $\mathcal{L}^{\to}_{A^{n}B^{n}\to X_{A}A^{\prime}B^{\prime}}$ from $n$ copies of $\rho_{AB}$ is equal to zero. This conclusion holds for every one-way LOCC channel $\mathcal{L}^{\to}_{A^{n}B^{n}\to X_{A}A^{\prime}B^{\prime}}$, every super two-extendible state $\rho_{AB}$, and every $n\in\mathbb{N}$. Hence, the one-shot probabilistic one-way distillable secret key of $\left(\rho_{AB}\right)^{\otimes n}$ is equal to zero for all $n\in\mathbb{N}$, and consequently, the probabilistic one-way distillable entanglement of every super two-extendible state is equal to zero.
@@ Appendix G Proof of Proposition 2
$\omega_{ABE}\coloneqq p~{}\gamma^{k}_{AB}\otimes[e]_{E}+(1-p)\gamma^{k}_{AE}\otimes[e]_{B},$ (91)
@@
$\operatorname{Tr}_{E}\!\left[\omega_{ABE}\right]$ $=p~{}\gamma^{k}_{AB}+(1-p)\gamma^{k}_{A}\otimes[e]_{B}=\widetilde{\eta}^{p,k}_{AB},$ (92)
$\operatorname{Tr}_{B}\!\left[\omega_{ABE}\right]$ $=(1-p)\gamma^{k}_{AE}+p~{}\gamma^{k}_{A}\otimes[e]_{E}=\widetilde{\eta}^{1-p,k}_{AE},$ (93)
@@
$E^{u}_{\min}\!\left(\widetilde{\eta}^{p,k}_{AB}\right)\leq\frac{1}{2}D_{\min}\!\left(\widetilde{\eta}^{p,k}_{AB}\Big{\|}\widetilde{\eta}^{1-p,k}_{AB}\right)=0\quad\forall p\in(0,1),k\geq 2.$ (94)
@@
$E^{u}_{\min}\!\left(\widetilde{\eta}^{p,k}_{AB}\right)=0\quad\forall p\in(0,1),k\geq 2.$ (95)
Finally, the erased private state $\widetilde{\eta}^{p,k}_{AB}$ is a separable state for $p=0$, and the min-unextendible entanglement of a separable state is equal to zero [22, Proposition 3]. Thus, we conclude the statement of Proposition 2.
@@ Appendix I Coherent information of Werner states
$I\!\left(A\rangle B\right)_{W^{p,d}}\\ =1-h_{2}(p)-p\log_{2}\!\left({d+1}\right)-(1-p)\log_{2}\!\left({d-1}\right).$ (105)
@@
where $\Pi^{\operatorname{sym}}$ and $\Pi^{\operatorname{asym}}$ are the projections on the symmetric and asymmetric subspaces of the underlying Hilbert space, respectively. For a bipartite system, the symmetric and asymmetric projection operators can be written as follows:
@@
Now let us evaluate the von Neumann entropy of the Werner state $W^{p,d}_{AB}$,
@@
$H\!\left(W^{p,d}_{AB}\right)=-p\frac{2}{d(d+1)}\log_{2}\!\left(p\frac{2}{d(d+1)}\right)\operatorname{Tr}\!\left[\Pi^{\operatorname{sym}}_{AB}\right]\\ -(1-p)\frac{2}{d(d-1)}\log_{2}\!\left((1-p)\frac{2}{d(d-1)}\right)\operatorname{Tr}\!\left[\Pi^{\operatorname{asym}}_{AB}\right]$ (120)
@@
$H\!\left(W^{p,d}_{AB}\right)$ $=-p\log_{2}\!\left(p\frac{2}{d(d+1)}\right)$
$\quad\qquad-(1-p)\log_{2}\!\left((1-p)\frac{2}{d(d-1)}\right)$ (123)
$=h_{2}(p)+p\log_{2}(d(d+1))-p$
$\qquad\quad+(1-p)\log_{2}(d(d-1))-(1-p)$ (124)
$=\log_{2}d+h_{2}(p)-1+p\log_{2}(d+1)$
$\quad\qquad+(1-p)\log_{2}(d-1),$ (125)
@@
$H\!\left(\zeta^{F,d}_{AB}\right)$
$=-\operatorname{Tr}\!\left[\zeta^{F,d}_{AB}\log_{2}\zeta^{F,d}_{AB}\right]$ (130)
$=-\operatorname{Tr}\!\left[F\Phi^{d}_{AB}\log_{2}\!\left(F\Phi^{d}_{AB}\right)\right]$
$\quad-\operatorname{Tr}\!\left[\frac{1-F}{d^{2}-1}\left(I_{AB}-\Phi^{d}_{AB}\right)\log_{2}\!\left(\frac{1-F}{d^{2}-1}\left(I_{AB}-\Phi^{d}_{AB}\right)\right)\right]$ (131)
$=-F\log_{2}F\operatorname{Tr}\!\left[\Phi^{d}_{AB}\right]$
$\quad-\frac{1-F}{d^{2}-1}\log_{2}\!\left(\frac{1-F}{d^{2}-1}\right)\operatorname{Tr}\!\left[I_{AB}-\Phi^{d}_{AB}\right]$ (132)
@@
$=h_{2}(F)+(1-F)\log_{2}\!\left(d^{2}-1\right)$ (135)
where we have used the property of projection operators mentioned in (119) to arrive at the third equality. Substituting (129) and (135) in the definition of coherent information, we arrive at (127).

### v2

@@ Appendix D The set of super two-extendible states is convex but not closed
$\operatorname{supp}\!\left(\operatorname{id}_{E\to B}(\operatorname{Tr}_{B}\!\left[\overline{\rho}_{ABE}\right])\right)\subseteq\operatorname{supp}\!\left(\rho_{AB}\right),$ (70)
where both systems $E$ and $B$ are modeled by the same Hilbert space. Note that the existence of such an extension is guaranteed by the fact that $\rho_{AB}$ is in the set of super two-extendible states. Similarly, let $\overline{\tau}_{ABE}$ be an extension of $\tau_{AB}$ such that the following containment holds:
$\operatorname{supp}\!\left(\operatorname{id}_{E\to B}(\operatorname{Tr}_{B}\!\left[\overline{\tau}_{ABE}\right])\right)\subseteq\operatorname{supp}\!\left(\tau_{AB}\right).$ (71)
From now on, we will assume that the identity channel $\operatorname{id}_{E\to B}$ is implicit when we compare the supports of two operators in $\mathcal{S}(AB)$ and $\mathcal{S}(AE)$, respectively.
@@
$\omega_{ABE}\coloneqq p\penalty 10000\ \overline{\rho}_{ABE}+(1-p)\overline{\tau}_{ABE},$ (73)
@@
$\operatorname{Tr}_{E}\!\left[\omega_{ABE}\right]=p\penalty 10000\ \rho_{AB}+(1-p)\tau_{AB}=\sigma_{AB}.$ (74)
@@
$\operatorname{supp}\!\left(\operatorname{Tr}_{B}\!\left[\omega_{ABE}\right]\right)$ $=\operatorname{supp}\!\left(\operatorname{Tr}_{B}\!\left[\overline{\rho}_{ABE}\right]\right)\cup\operatorname{supp}\!\left(\operatorname{Tr}_{B}\!\left[\overline{\tau}_{ABE}\right]\right)$ (76)
@@
Since the support of $\operatorname{Tr}_{B}\!\left[\omega_{ABE}\right]$ is in the support of $\operatorname{Tr}_{E}\!\left[\omega_{ABE}\right]$, we conclude that $\operatorname{Tr}_{E}\!\left[\omega_{ABE}\right]=\sigma_{AB}$ is a super two-extendible state for all $p\in(0,1)$ as well. Therefore, the state $\sigma_{AB}$, which is a convex combination of two arbitrary super two-extendible states, is super two-extendible for every $p\in[0,1]$, justifying that the set of super two-extendible states is convex.
To show that the set of super two-extendible states is not closed, let us consider the example of erased states defined in (16). The erased state $\widetilde{\Phi}^{p,k}_{AB}$ is a special case of an erased private state $\widetilde{\eta}^{p,k}_{AB}$. Appendix G establishes that an erased private state is super two-extendible for every $p\in[0,1)$. This implies that an erased state $\widetilde{\Phi}^{p,k}_{AB}$ is also super two-extendible for every $p\in[0,1)$. However, for $p=1$, $\widetilde{\Phi}^{p,k}_{AB}$ is a maximally entangled state of Schmidt rank $k$, which is not super two-extendible. In fact, this state has a min-unextendible entanglement of $\log_{2}k$. As such, the maximally entangled state is a limit point of the set of super two-extendible states, but it is not a super two-extendible state itself. Hence, the set of super two-extendible states is not closed.
@@ Appendix E Proof of Proposition 1
We now establish the opposite implication. Consider an arbitrary quantum state $\rho_{AB}$ such that $E^{u}_{\min}\!\left(\rho_{AB}\right)=0$. This is true only if there exists a quantum state $\sigma_{AB}\in\mathcal{F}\!\left(\rho_{AB}\right)$ such that $\operatorname{Tr}\!\left[\Pi^{\rho}_{AB}\sigma_{AB}\right]=1$, which follows from the definition of the min-unextendible entanglement of a bipartite state. From the gentle measurement lemma [38, Lemma 5], it is known for a state $\tau$ and a projector $P$ that
@@ Appendix F Proof of Theorem 1
$0$ $\leq E^{u}_{\min}\!\left(q[1]_{X_{A}}\otimes\gamma^{k^{\prime}}_{A^{\prime}B^{\prime}}+(1-q)[0]_{X_{A}}\otimes\frac{I_{A^{\prime}B^{\prime}}}{d_{A^{\prime}}d_{B^{\prime}}}\right)$ (88)
@@
Inspecting (88) and (93), it follows that $E^{u}_{\min}\!\left(\eta^{q,k^{\prime}}_{A^{\prime}B^{\prime}}\right)=0$. By applying Lemma 1, using the fact that $\frac{1}{2}\log_{2}\!\left(\frac{q}{k^{\prime 2}}+1-q\right)\geq 0$ for all $q\in[0,1]$ and integer $k^{\prime}\geq 2$, and solving the equation $0=-\frac{1}{2}\log_{2}\!\left(\frac{q}{k^{\prime 2}}+1-q\right)$, along with the assumption that $k^{\prime}\geq 2$, we conclude that $q=0$. As such, the expected number of secret key bits distilled by the channel $\mathcal{L}^{\to}_{A^{n}B^{n}\to X_{A}A^{\prime}B^{\prime}}$ from $n$ copies of $\rho_{AB}$ is equal to zero. This conclusion holds for every one-way LOCC channel $\mathcal{L}^{\to}_{A^{n}B^{n}\to X_{A}A^{\prime}B^{\prime}}$, every super two-extendible state $\rho_{AB}$, and every $n\in\mathbb{N}$. Hence, the one-shot probabilistic one-way distillable secret key of $\left(\rho_{AB}\right)^{\otimes n}$ is equal to zero for all $n\in\mathbb{N}$, and consequently, the probabilistic one-way distillable secret key of every super two-extendible state is equal to zero.
@@ Appendix G Proof of Proposition 2
$\omega_{ABE}\coloneqq p\penalty 10000\ \gamma^{k}_{AB}\otimes[e]_{E}+(1-p)\gamma^{k}_{AE}\otimes[e]_{B},$ (94)
@@
$\operatorname{Tr}_{E}\!\left[\omega_{ABE}\right]$ $=p\penalty 10000\ \gamma^{k}_{AB}+(1-p)\gamma^{k}_{A}\otimes[e]_{B}=\widetilde{\eta}^{p,k}_{AB},$ (95)
$\operatorname{Tr}_{B}\!\left[\omega_{ABE}\right]$ $=(1-p)\gamma^{k}_{AE}+p\penalty 10000\ \gamma^{k}_{A}\otimes[e]_{E}=\widetilde{\eta}^{1-p,k}_{AE},$ (96)
@@
$E^{u}_{\min}\!\left(\widetilde{\eta}^{p,k}_{AB}\right)\leq\frac{1}{2}D_{\min}\!\left(\widetilde{\eta}^{p,k}_{AB}\Big\|\widetilde{\eta}^{1-p,k}_{AB}\right)=0\qquad\forall p\in(0,1),k\geq 2.$ (97)
@@
$E^{u}_{\min}\!\left(\widetilde{\eta}^{p,k}_{AB}\right)=0\qquad\forall p\in(0,1),k\geq 2.$ (98)
Finally, the erased private state $\widetilde{\eta}^{p,k}_{AB}$ is a separable state for $p=0$, and the min-unextendible entanglement of a separable state is equal to zero [25, Proposition 3]. Thus, we conclude the statement of Proposition 2.
@@ Appendix I Coherent information of Werner states
$I\!\left(A\rangle B\right)_{W^{p,d}}=1-h_{2}(p)-p\log_{2}\!\left({d+1}\right)-(1-p)\log_{2}\!\left({d-1}\right).$ (108)
@@
where $\Pi^{\operatorname{sym}}$ and $\Pi^{\operatorname{asym}}$ are the projections onto the symmetric and asymmetric subspaces of the underlying Hilbert space, respectively. This is easily seen from the fact that the symmetric and asymmetric projection operators for a bipartite system can be written as follows:
@@
Now let us evaluate the von Neumann entropy of the Werner state $W^{p,d}_{AB}$:
@@
$H\!\left(W^{p,d}_{AB}\right)=-p\frac{2}{d(d+1)}\log_{2}\!\left(p\frac{2}{d(d+1)}\right)\operatorname{Tr}\!\left[\Pi^{\operatorname{sym}}_{AB}\right]\\ -(1-p)\frac{2}{d(d-1)}\log_{2}\!\left((1-p)\frac{2}{d(d-1)}\right)\operatorname{Tr}\!\left[\Pi^{\operatorname{asym}}_{AB}\right].$ (123)
@@
$H\!\left(W^{p,d}_{AB}\right)$ $=-p\log_{2}\!\left(p\frac{2}{d(d+1)}\right)-(1-p)\log_{2}\!\left((1-p)\frac{2}{d(d-1)}\right)$ (126)
$=h_{2}(p)+p\log_{2}(d(d+1))-p+(1-p)\log_{2}(d(d-1))-(1-p)$ (127)
$=\log_{2}d+h_{2}(p)-1+p\log_{2}(d+1)+(1-p)\log_{2}(d-1),$ (128)
@@
$H\!\left(\zeta^{F,d}_{AB}\right)$ $=-\operatorname{Tr}\!\left[\zeta^{F,d}_{AB}\log_{2}\zeta^{F,d}_{AB}\right]$ (133)
$=-\operatorname{Tr}\!\left[F\Phi^{d}_{AB}\log_{2}\!\left(F\Phi^{d}_{AB}\right)\right]-\operatorname{Tr}\!\left[\frac{1-F}{d^{2}-1}\left(I_{AB}-\Phi^{d}_{AB}\right)\log_{2}\!\left(\frac{1-F}{d^{2}-1}\left(I_{AB}-\Phi^{d}_{AB}\right)\right)\right]$ (134)
$=-F\log_{2}F\operatorname{Tr}\!\left[\Phi^{d}_{AB}\right]-\frac{1-F}{d^{2}-1}\log_{2}\!\left(\frac{1-F}{d^{2}-1}\right)\operatorname{Tr}\!\left[I_{AB}-\Phi^{d}_{AB}\right]$ (135)
@@
$=h_{2}(F)+(1-F)\log_{2}\!\left(d^{2}-1\right),$ (138)
where we have used the property of projection operators mentioned in (122) to arrive at the third equality. Substituting (132) and (138) in the definition of coherent information, we arrive at (130). This concludes the proof.